OpenStudy (anonymous):

Solve: 4x-3y=-2 -3x + 2y =0

OpenStudy (anonymous):

u can try adding the system of equations, 4x-3y=-2 + -3x+2y=0 = x-y=-2 bring the y over to the other side, x=y-2 plug that in for the x values and solve 4(y-2)-3y=-2 4y-8-3y=-2 y-8=-2 (add 8 on both sides) y=6 plug y in and then do the same for x

OpenStudy (anonymous):

Write one of them in the form y= ax + c So: -3x+2y = 0 becomes y=1.5x Replace the y in 4x-3y = -2 with 1.5x ---> 4x -3(1.5x) = -2 4x-4.5x=-2 -0.5x=-2 x=4 Then plug x=4 in the equation: -3*4+2y=0 2y=12 y=6 So: (4,6) is the intersect point.

OpenStudy (anonymous):

so then 4x-3(6)=-2 4x-18=-2 4x=16 x=4